OpenStudy (anonymous):

2/3 / 8/3 + -7/6 X 1/6

OpenStudy (anonymous):

if you divide the fractions, use the reciprocal of the second and multiply.

OpenStudy (anonymous):

i tried =/

OpenStudy (anonymous):

so you get 2/3 * 3/8=2/8=1/4

OpenStudy (anonymous):

i got that

OpenStudy (anonymous):

its wrong

OpenStudy (anonymous):

do the same with the second part

OpenStudy (anonymous):

i got 1/18 its right

OpenStudy (anonymous):

i'll show it in a drawing, give me a minute...

OpenStudy (anonymous):